### v2

# Multi-AUV Cooperative Underwater Multi-Target Tracking Based on Dynamic-Switching-enabled Multi-Agent Reinforcement Learning

Abstract
With the rapid development of underwater communication, sensing, automation, robot technologies, autonomous underwater vehicle (AUV) swarms are gradually becoming popular and have been widely promoted in ocean exploration and underwater tracking or surveillance, etc. However, the complex underwater environment poses significant challenges for AUV swarm-based accurate tracking for the underwater moving targets. In this paper, we aim at proposing a multi-AUV cooperative underwater multi-target tracking algorithm especially when the real underwater factors are taken into account. We first give normally modelling approach for the underwater sonar-based detection and the ocean current interference on the target tracking process. Then, we regard the AUV swarm as a underwater ad-hoc network and propose a novel Multi-Agent Reinforcement Learning (MARL) architecture towards the AUV swarm based on Software-Defined Networking (SDN). It enhances the flexibility and scalability of the AUV swarm through centralized management and distributed operations. Based on the proposed MARL architecture, we propose the ”dynamic-attention switching” and ”dynamic-resampling switching” mechanisms, to enhance the efficiency and accuracy of AUV swarm cooperation during task execution. Finally, based on a proposed AUV classification method, we propose an efficient cooperative tracking algorithm called ASMA. Evaluation results demonstrate that our proposed tracking algorithm can perform precise underwater multi-target tracking, comparing with many of recent research products in terms of convergence speed and tracking accuracy.
Index Terms:
AUV swarm network, multi-target tracking, dynamic-switching attention, dynamic-switching resampling, software-defined networking.
I Introduction
Covering more than 70% of Earth’s surface, the ocean is a vital life-support system with abundant biological and mineral resources [1, 2, 3]. With the advancement of technology and our increasing knowledge of the oceans, the need for efficient and intelligent exploration technologies, particularly in ocean environment surveillance [4, 5], disaster prevention, and rescue operations [6, 7], becomes increasingly pressing. In these years, the rapid development of underwater robots, underwater communication, underwater sensing technologies, has given birth to the Autonomous Underwater Vehicle (AUV). As of today, AUV has become indispensable tools in a wide range of marine applications especially in underwater target tracking [8, 9], due to its ability to autonomously navigate underwater environments.
With the continuous improvement and development of swarm intelligence theory, the control technology of AUV swarms has been greatly propelled and enhanced [10, 11]. The AUV swarms have demonstrated unique advantages in executing tasks in complex marine environments, e.g., the AUV-swarm-based cooperative underwater data collection [12], cooperative underwater target tracking [13]. Through cooperation manner, the AUVs in AUV swarm can cover broader areas, enhancing the efficiency and accuracy of underwater operatiion. Particularly in underwater cooperative multi-target tracking, the AUV swarms have been demonstrated exceptional capabilities in cooperatively tracking various underwater targets such as underwater biological communities [14] or environmental pollutants [15]. This is crucial not only for underwater scientific research such as ecological [16] surveys but also for various categories of civil [17] and military applications [18], such as supporting maritime search, rescue operations and maritime boundary surveillance, etc.
However, effectively managing and cooperating such a complex multi-agent system, especially in dynamic and unpredictable marine environments, has to face many challenges. Particularly in underwater multi-target tracking, these challenges mainly manifest in several aspects: The complexity and dynamism of the marine environment, such as ocean currents interference [19, 20], this demands AUV swarms to possess high adaptability and flexibility to cope with sudden environmental changes and unknown underwater obstacles; Effective communication is crucial for guaranteeing the AUV swarm-based underwater cooperative operations, but the high latency and low bandwidth characteristics of underwater communication especially underwater acoustic-based communication limit the efficiency of the information exchange among the AUVs in AUV swarm and cannot support flexible swarm cooperation [21, 22, 23]. Thus, it requires a scalable network architecture to re-organize and re-define the information exchange architecture for AUV swarm system; Meanwhile, in reality, due to productivity constraints, AUV swarms often comprise various heterogeneous AUVs, making AUV swarm-based cooperative underwater operation more challenging and impeding support for cooperative swarm intelligence [24]. Further, traditional MARL algorithms suffer from long runtimes, slow convergence speeds, and high resource consumption, making MARL challenging to deploy in large-scale AUV swarm.
As a cutting-edge networking approach, Software-Defined Networking (SDN) separates the control plane from the data plane in traditional network architectures [25, 26, 27, 28]. By this architecture, the network control functions are centralized into one or more controllers. This allows the network resources together with the network information to be flexibly managed and utilized. Motivated by this, in this work, we display how to utilize SDN technology to re-define the AUV swarm as a AUV swarm network with self-learning ability. For dedicated MARL architecture towards AUV swarm network, we propose a MARL algorithm incorporating a dynamic-switching mechanism, which demonstrates faster convergence and better performance in multi-target tracking tasks under ocean current interference. Totally, this paper mainly makes the following contributions:
  1. 1.
    We regard the AUV swarm as a AUV swarm network, utilize SDN to optimize AUV swarm network and enhance the flexibility and scalability of the AUV swarm;
  2. 2.
    We propose a multi-agent reinforcement learning algorithm incorporating dynamic-switching mechanism (including both ”dynamic-attention switching” and ”dynamic-resampling switching”) to enhance AUV swarm network’s self-learning efficiency and accuracy;
  3. 3.
    We introduces an AUV formation classification algorithm based on fuzzy logic and rule-based expert systems, to enable the AUV swarm network to perform precise tracking of multiple targets in the interference of ocean currents.
The rest of this paper is organized as follows: Related works on the paper’s theme is surveyed in Sec. II; Preliminary materials are introduced in Sec. III; The proposed FATHOM-Net is detailed in Sec. IV; The proposed cooperative tracking algorithm is presented in Sec. V; Evaluation results are shown in Sec. VI; The paper concludes in Sec. VII with a discussion of some future research directions derived from this paper.
II Related Works
In this section, we investigate recent advances related to the main research topics, specifically categorized into the following three areas: 1) SDN-based underwater multi-agent system, 2) Single AUV-based target tracking, and 3) AUV swarm-based target tracking.
II-A SDN-based underwater multi-agent system
SDN technology, characterized by its network architecture that separates the control plane from the data forwarding plane, provides strong support for centralized control and flexible management of underwater multi-agent systems. This technology is currently widely applied in fields such as underwater multi-agent networks, opening new possibilities for underwater operations and research.
For instance, in [29], the authors propose the DRSIR model, a deep reinforcement learning approach for routing selection in SDN. This method utilizes path state indicators and demonstrates higher efficiency and intelligence compared to traditional Dijkstra and RSIR algorithms. It can dynamically adjust routing strategies based on changes in network traffic, showcasing practical feasibility and superior performance in SDN routing.
In [30], an SDN-based architecture for AUV underwater wireless networks is introduced to support multi-AUV cooperative search. This architecture achieves network information synchronization, node localization, multi-AUV cooperative control, and intelligent data transmission scheduling through software-defined beaconing, hierarchical localization, cooperative control, and software-defined hybrid data transmission frameworks. Building upon these studies, we leverage SDN technology to enhance the underwater vehicle swarm network system.
II-B Single AUV-based target tracking
In Single AUV-based target tracking, the focus is on enhancing the perception [31], decision-making [32], and control capabilities [33] of an individual AUV to autonomously perform identification and tracking of specific targets. This approach necessitates highly integrated sensor systems and advanced data processing algorithms to adapt to complex underwater environments and achieve high-precision tracking.
In [33], the authors develop a DRL-based control strategy for X-rudder AUVs, utilizing the DDPG algorithm to enable precise posture control and efficient target tracking. This approach achieves precise three-degree-of-freedom posture control and rapid DRL algorithm deployment, and significantly enhances AUV maneuverability for dynamic underwater target tracking.
In [32], the authors introduce an optimized AUV path planning method (PPM-BBD) employing L-SHADE, focusing on energy-efficient diving through decision-making under motion constraints. This approach minimizes energy consumption by up to 9% compared to conventional methods, demonstrating the importance of strategic decision-making in solitary AUV target tracking amidst environmental challenges.
II-C AUV swarm-based target tracking
AUV swarm-based target tracking leverages the coordinated efforts of multiple AUVs, utilizing principles of swarm intelligence to improve tracking range, efficiency, and robustness. This method not only extends beyond the limitations of a single AUV but also emphasizes distributed decision-making and information sharing among the AUVs to effectively track multiple or wide-area targets in dynamic environments [34, 35].
In [34], the authors propose a novel multi-AUV cooperative tracking scheme named CTDE, which employs a central training and distributed execution approach to achieve safe and effective tracking of moving targets using MADDPG. By conducting central training within a designed secure private network, information sharing is not required during task execution, thereby enhancing the security of the entire system.
In [35], the authors discuss trajectory tracking control of multiple AUVs using discrete-time control methods in weak communication environments. The study focuses on considering communication delays (bounded and unbounded) and packet loss scenarios, ensuring stable tracking of given trajectories by designing methods based on leader-follower and virtual leader approaches.
Totally, multi-AUV-based underwater target tracking technology has received widespread research attention. However, this research field still faces several challenges, such as adaptability to dynamic and complex marine environments, flexibility and robustness of AUV swarm networks, and how to stably and accurately track multiple targets. In response to these issues, our work employ the SDN technology to improve the scalability and functionability of the AUV swarm system. On account of the above system architecture, we propose a dynamic-switching-enabled MARL architecture which is more feasible for AUV swarm system’s self-learning. And, we propose as well a novel AUV classification method to improve the scalability of the AUV swarm-based tracking system. In the following, we will prove our work demonstrates precise multi-target tracking in dynamic and complex marine environments, compared to many existing studies.
III Preliminary Materials
In order to accurately track underwater targets and adapt to the dynamic factors in real underwater environments, we consider the effects of sonar detection and ocean current interference. Furthermore, we transform the target tracking scenario based on AUV swarm network into a Markov Decision Process (MDP).
III-A Sonar detection modeling
In real complex underwater environments, traditional electromagnetic-based detection methods are limited in their effectiveness. Therefore, in this paper, we utilize sonar technology to detect AUVs’ position together with their tracking targets. Equipped with sonar, AUVs emit directional sound waves to scan their surrounding environment. This process involves using a sector array sonar to capture echo signals from different directions, and the variations in signal intensity allow us to estimate the position of targets. In this study, a specific active sonar equation model is employed to simulate and optimize the target detection process, which can be expressed by Eq. 1.
$EM=SL-2TL+TS-(NL-DI)-DT$ (1)
where $SL$ is the strength of the sonar signal emitted, $TL$ is the attenuation of sound propagation, $TS$ is the target echo strength, $NL$ is the ambient marine noise, $DI$ is the system’s noise suppression capability, and $DT$ is the minimum signal-to-noise ratio required for the equipment to function properly. The units for the notations mentioned are all in decibels (dB).
III-B Ocean current modeling
The underwater environments’ uncertainty, especially the influence of ocean currents, poses significant challenges to the navigation and positioning capabilities of AUVs. It demands AUVs to possess a high level of adaptability and intelligence.
The Navier-Stokes equations are the fundamental equations for describing fluid dynamics behavior. They can simulate the characteristics of fluid flow. Additionally, they provide a computational environment for calculating the forces exerted by ocean currents (Eq. 2).
$\rho\left(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}\right)=-\nabla p+\mu\nabla^{2}\mathbf{u}+\mathbf{F}$ (2)
where $\rho$ is fluid density, $\mathbf{u}$ is velocity, $\frac{\partial\mathbf{u}}{\partial t}$ is velocity change over time, $\mathbf{u}\cdot\nabla\mathbf{u}$ is self-advection, $\nabla p$ is pressure change, $\mu$ is viscosity, $\nabla^{2}\mathbf{u}$ is the diffusion term, and $\mathbf{F}$ is external forces.
However, the Navier-Stokes equations involve a large number of partial differential calculations, resulting in high computational costs and long processing times. To reduce the complexity of the model and obtain preliminary fluid dynamic characteristics, we assume that fluid flow is laminar and relatively uniform. Neglecting the nonlinear convective term $\mathbf{u}\cdot\nabla\mathbf{u}$ in the Navier-Stokes equations simplifies it to a linear problem that can be expressed by the following.
$\rho\frac{\partial\mathbf{u}}{\partial t}=-\nabla p+\mu\nabla^{2}\mathbf{u}+\mathbf{F}$ (3)
After calculating $\mathbf{u}$ and $\frac{\partial\mathbf{u}}{\partial t}$ through the Navier-Stokes equations, we use the drag equation (Eq. 4), lift equation (Eq. 5), and virtual mass equation (Eq. 6) to calculate the forces acting on the structure under the influence of ocean currents:
$F_{D}=\frac{1}{2}\rho u^{2}C_{D}A$ (4)
where $F_{D}$ is the drag force, $\rho$ is fluid density, $u$ is relative fluid velocity, $C_{D}$ is drag coefficient, and $A$ is object’s frontal area.
$F_{L}=\frac{1}{2}\rho u^{2}C_{L}A$ (5)
where $F_{L}$ is the lift force, $C_{L}$ is lift coefficient.
$F_{VM}=\rho C_{VM}V\frac{\partial\mathbf{u}}{\partial t}$ (6)
where $F_{VM}$ is the virtual mass force, $C_{VM}$ is the virtual mass coefficient, $V$ is AUV’s volume, and $\frac{\partial u}{\partial t}$ is the rate of change of velocity.
Through the above equations (Eq. 4-Eq. 6), we can compute the complex and variable forces exerted by ocean currents on the AUV. The three forces are shown in Fig. 1. The overall calculation for the ocean current can be summarized in Algorithm 1.
Figure 1: Schematic diagram of ocean current
Algorithm 1 Simplified fluid dynamics calculation algorithm
1:Initialize the Set $N_{E}$, $T$ respectively.
2:Initialize AUV’s frontal area $A$, volume V, drag coefficient $C_{D}$, lift coefficient $C_{L}$, and inertia coefficient $C_{VM}$.
3:for $\text{Episode}=1$ to $N_{E}$ do
4:     Initialize fluid density $\rho$, dynamic viscosity $\mu$, and external forces $\mathbf{F}$.
5:     Initialize environment and fluid velocity $\mathbf{u}$ and pressure gradient $\nabla p$.
6:     for $t=1$ to $T$ do
7:         Calculate the velocity time derivative $\frac{\partial\mathbf{u}}{\partial t}$ by simplified Navier-Stokes equation.
8:         Update fluid velocity $\mathbf{u}$ and pressure gradient $\nabla p$.
9:         Calculate the drag force $F_{D}$, lift force $F_{L}$, and virtual mass force $F_{VM}$ by Eq. 4 to Eq. 6.
10:         Combine forces to determine net force on the AUV.
III-C Markov decision process modeling
In the decision-making process of AUV swarm network, the environment is defined as a MDP (Eq. 7), consisting of five main components: the state set $\mathcal{S}$, the action set $\mathcal{A}$, the state transition probabilities $\mathcal{P}$, the reward function $\mathcal{R}$, and the discount factor $\gamma$:
$\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$ (7)
The state set $\mathcal{S}$ encompasses the sonar detection results in different coordinate systems and their location information. Specifically, it is a set composed of the target and the AUV’s echo intensity along with their location information. $\mathcal{S}$ can be represented as $\mathcal{S}=\{s_{1},\ldots,s_{N}\}$, where $s_{i}=(\eta_{i},\phi_{i},o_{i})$, $o_{i}=(\kappa_{i},\sigma_{i})\in\mathbb{R}^{N_{\kappa}+N_{\sigma}}$, $\kappa_{i}$ represents the echo intensity from the target, $\sigma_{i}$ represents the echo intensity from other AUVs, $N_{\kappa}$ is the number of targets being tracked, and $N_{\sigma}$ is the number of other AUVs. The action set $\mathcal{A}$ in this study represents the discrete action set of all AUVs, denoted as $\mathcal{A}=\{a_{1},\ldots,a_{N}\}$, where each action $a_{i}$ is a vector representing seven possible actions (e.g., up, down, left, right, forward, backward, and stay in place). The state transition probabilities $\mathcal{P}$ describe the dynamic transition relationship between states, which is key for the algorithm to understand how different states transition into each other. The discount factor $\gamma$, where the value of $\gamma$ ranges between 0 and 1, together determines the calculation method for long-term benefits in policy evaluation. Further, the reward function $\mathcal{R}$ will be explained in Sec. V, where we consider tracking accuracy, collision avoidance, ocean current stability, and energy efficiency to enhance multi-AUV cooperative tracking effects.
IV Hierarchical Software-Defined Multi-AUV Reinforcement Learning Architecture
In this section, we introduce Flexible Autonomous Training Hierarchy for Operational Marine Network (FATHOM-Net), which is a software-defined MARL architecture based on Centralized Training and Decentralized Execution.
Figure 2: Proposed FATHOM-Net
IV-A Overview of FATHOM-Net
FATHOM-Net framework consists of three functional layers: global-control training layer, local-control training layer, and application execution tracking layer, which can be detailed as the following.
Global-control training layer: The global-control training layer comprises a global controller based on Unmanned Surface Vessel-based Global Controller (USV-GC). It aims to coordinate Local Controller-based AUV (LC-AUV)s in various regions through a dedicated northbound interface, ensuring efficient classification of large-scale tasks to suitable local-control training layers by deconstructing and optimizing tasks. The strategy updates and task scheduling mechanism in this layer not only support the efficient execution of complex tasks but also continuously optimize the global network view by periodically obtaining status information from the lower layers. This ensures the flexibility and efficiency of network operations, especially when facing environmental challenges and changing task requirements.
Local-control training layer: It focuses on optimizing tasks issued by the global-control training layer into specific strategies through training, and conveying them to each managed Execution Tracking-based AUV (ET-AUV). Upon assignment of subregion tasks by the global-control training layer, each LC-AUV begins centralized local training. This process uses the unique characteristics of the ET-AUV subset it manages, following the parameters and tasks provided by the global-control training layer. This further refines tasks and formulates specific execution strategies to guide its managed ET-AUVs. Additionally, it periodically maintains the local network view through a dedicated southbound interface. As LC-AUVs require a certain amount of computation, they are typically equipped with high-performance batteries and computing units. This training mode ensures that all ET-AUVs can execute tasks according to uniform standards and optimal strategies, while also allowing for rapid adjustments tailored to local environments and specific task requirements.
Application execution tracking layer: It is directly responsible for executing specific strategies. This layer consists of ordinary ET-AUVs swarms. Each ET-AUV independently executes tasks and strategies received from the local-control training layer. Therefore, communication among ET-AUVs is not imposed. This idea (the idea of ”separate control from operation” in SDN) reduces communication costs and enhances the flexibility and robustness of the AUV swarm network. Even if individual ET-AUVs fails or are destroyed, it does not affect the overall efficiency of the network.
Overall, FATHOM-Net reduces the coupling between individual AUVs, addressing the scalability issues present in traditional architectures. The FATHOM-Net framework is summarized in Fig. 2. The entire task classification processing is summarized in Fig. 3.
IV-B Beacon framework for FATHOM-Net
In order to ensure precise target tracking based on the proposed FATHOM-Net, it is crucial to establish both local and global network views for LC-AUVs and USV-GC. Therefore, we propose an SDN beacon framework to synchronize and share information among ET-AUVs, LC-AUVs, and USV-GC. The proposed SDN beacon framework can be carried out by the following:
Phase 1: To update local and global network views, the beaconing is activated between USV-GC/LC-AUVs and their subordinate levels within the region requesting synchronization (SYN) information. Upon receiving a synchronization request, each LC-AUV/ET-AUV replies with key information such as its AUV ID and status to update the network views.
Phase 2: To allocate specific-tasks/strategies to LC-AUVs/ET-AUVs through centralized training. In each training round, USV-GC/LC-AUVs send requests to LC-AUVs/ET-AUVs they manage after training completion, and send specific-tasks/strategies to them through operation requests. Each LC-AUV/ET-AUV then replies with key information.
In summary, Phase 1 describes the network view updating procedure, while Phase 2 displays the task classification processing. To improve the information exchange efficiency, the reply information in Phase 2 also includes information for updating the network views. Therefore, these replies are also considered as an update to the network views.
V Proposed Cooperative Tracking Algorithm
In this section, we showcase our proposed cooperative tracking algorithm. Especially, we firstly propose a classification algorithm called AUV Scalable Management and Allocation (ASMA) to classify/allocate ET-AUVs to each formation for performing cooperative multi-target tracking. To enhance the tracking accuracy and the convergence speed of the AUV swarm network, we employ the concept of ”Dynamic-Switching”, integrate the ”Dynamic-Switching Attention” mechanism into the Actor-Critic model and develop the Dymiac-Switching-Based MARL (DSBM) architecture. Additionally, to improve the efficiency of better sample utilization in DSBM, we introduce a novel ”Dynamic-Switching Resampling” mechanism based on experience replay buffer mechanism.
V-A AUV classification method
In underwater multi-target tracking problems, efficiently and reasonably classifying tracking formations (consisting of a series of ET-AUVs) for each underwater target is an important issue. Totally, the proposed ASMA achieves efficient and accurate formation assignments for ET-AUVs by utilizing a combination of fuzzy logic and rule-based expert systems, integrating various performance indicators.
The evaluation process initially employs fuzzy logic scoring to map specific performance indicators of ET-AUVs onto a membership degree. The mathematical expression for this function is as follows (Eq. 8):
Figure 3: Task classification based on FATHOM-Net
TABLE I: Special notations in Sec. V
Notation Description
$\alpha_{i}$ Acceleration score
$\beta_{i}$ Remaining battery score
$\kappa_{i}$ Carrying capacity score
$\epsilon_{i}$ Energy consumption score
$d_{i,j}$ Distance between ET-AUV $i$ to the target $j$
$D_{i}$ Perception range
$v_{i}$ Speed score
$\mu(x)=\begin{cases}0&\text{if }x<\text{range\_min}\\ 1&\text{if }x>\text{range\_max}\\ \frac{x-\text{range\_min}}{\text{range\_max}-\text{range\_min}}&\text{otherwise}\end{cases}$ (8)
where $\mu(x)$ is the membership degree of notation $x$ normalized between 0 and 1, $x$ is the actual value of the notation, and range_min and range_max set the normalization range for $x$, specifying the minimum and maximum values, respectively.
The rule-based expert system takes the fuzzy membership degrees as input. It then calculates the overall score $\mathit{S}_{i,j}$ acquired by agent $i$ and target $j$. The notations and their meanings are shown in Table I.
Thus, based on ASMA, the score $\mathit{S}_{i,j}$ acquired by ET-AUV $i$ and underwater tracking target $j$ can be computed by the following steps:
Step 1: Energy regulation score:
$\mathit{S}_{i,j}+=\left(v_{i}^{3}+2\alpha_{i}^{2}\right)\times\frac{1}{1+e^{-\beta_{i}}}\times\left(1-\frac{d_{i,j}}{D_{i}}\right)$ (9)
Significant increases in total $\mathit{S}_{i,j}$ occur when ET-AUVs have fast speeds, stable accelerations, and sufficient power. However, when the remaining power is low, even if the ET-AUV accelerates quickly, the $\mathit{S}_{i,j}$ will be lower to avoid that ET-AUVs quickly deplete their power. Additionally, as the distance between the ET-AUV and the target approaches the maximum perception range, the score decreases to prevent losing track of the target.
Step 2: Efficiency load balancing score:
$\mathit{S}_{i,j}+=\cosh(\epsilon_{i}-\kappa_{i})\times\sqrt{\lvert\kappa_{i}-\epsilon_{i}\rvert}$ (10)
By comparing energy consumption with payload capacity, the balance between energy efficiency and stability of ET-AUVs during task execution is emphasized.
Step 3: Comprehensive performance score:
$\mathit{S}_{i,j}+=\frac{v_{i}^{2}+\alpha_{i}^{2}+\epsilon_{i}^{2}}{3}$ (11)
This rule considers the comprehensive impact of speed, acceleration, and energy consumption, reflecting the stability of ET-AUV’s overall performance under different operating conditions.
Step 4: Dynamic counter-reaction score:
$\mathit{S}_{i,j}+=\frac{1}{1+e^{\alpha_{i}-v_{i}}}$ (12)
If the ET-AUV can smoothly adjust its speed, it indicates high-performance control capability under dynamic conditions. This is crucial for EC-AUVs to perform complex underwater evasive action, e.g., collision avoidance, precise positioning, efficient navigation, etc.
Step 5: Fine operational expertise score:
$\mathit{S}_{i,j}+=\text{if }v\leq 0.3\text{ then }v\times\kappa_{i}\text{ else}0$ (13)
In underwater tasks requiring precise control, ET-AUVs with low speed and high payload capacity will receive additional $\mathit{S}_{i,j}$ due to higher operational accuracy.
Step 6: Remote rapid response score:
$\mathit{S}_{i,j}+=\text{if }v\geq 0.7\text{ then }v\times\beta_{i}\times\frac{d_{i,j}}{D_{i}}\text{ else }0$ (14)
In underwater urgent tasks requiring the ET-AUV to fleetly arrive at the distant targets, high-speed ET-AUVs at greater distances from the target receive additional $\mathit{S}_{i,j}$, reflecting their rapid response capability to distant targets.
Totally, the proposed ASMA is carried out based on the $\mathit{S}_{i,j}$ of each ET-AUV for each target. This ensures that every ET-AUV tracks a target and each target is tracked by at least one ET-AUV. The optimization goal is to maximize the cumulative score.
In summary, ASMA can be summarized as Algorithm 2, which considers various special factors in different scenarios, classifying rational and efficient tracking formations.
Algorithm 2 ET-AUV performance classification algorithm
1:Initialize the number $k$ of ET-AUVs and the number $T$ of targets.
2:for $i=1$ to $k$ do
3:     Initialize ET-AUV $i$’s $\alpha_{i}$, $\beta_{i}$, $\kappa_{i}$, $\epsilon_{i}$, $d_{i,j}$, $D_{i}$ and $v_{i}$.
4:for $i=1$ to $k$ do
5:     Calculate membership values for ET-AUV $i$ by Eq. 8.
6:for $i=1$ to $k$ do
7:     for $j=1$ to $T$ do
8:         Update the score $\mathit{S}_{i,j}$ of ET-AUV $i$ for target $j$ by Eq. 9 to Eq. 14.
9:Assign the best formation plan based on score $\mathit{S}_{i,j}$.
V-B Proposed dynamic-switching-based MARL
Normally, in MARL, there are two major challenges: the high variance problem of traditional strategy estimation methods and the slow convergence problem caused by the high dimensionality of Critic networks in multi-agent systems. To address these issues, in this paper, we introduce Temporal Difference techniques to reduce evaluation variance and propose DSBM, which adopts the ”Dynamic-Switching” mechanism to improve learning efficiency and accelerate policy convergence. The ”Dynamic-Switching” mechanism focuses on optimizing the handling of high-value samples, as shown on the right side of Fig. 2.
Dynamic-switching attention mechanism: During the process of tracking targets, there are a total of $N$ ET-AUVs. For a specific ET-AUV $i$, we need to calculate its Q-value function $Q_{\psi_{i}}(o,a)$, where $o$ represents the current state and $a$ represents the action. The $Q_{\psi_{i}}(o,a)$, which represents the value of the action taken by ET-AUV $i$ given a particular observation and action. It can be obtained through the following steps.
For each ET-AUV, first identify the ET-AUV with the highest reward in the current training round ($AUVbe$) (Eq. 15).
$AUV_{\text{be}}=\text{argmax}(R_{f})$ (15)
where $R_{f}$ is the set of reward value for the AUVs at round $f$.
After that, it randomly select the experience from the remaining ET-AUVs (Eq. 16).
$AUV_{\text{ad}}=\begin{cases}AUV_{\text{be}}\cup\text{ran}(R^{*},2)&\text{if }i=AUV_{\text{be}}\\ AUV_{\text{be}}\cup\text{ran}(R^{*},1)&\text{if }i\neq AUV_{\text{be}}\end{cases}$ (16)
where $AUV_{\text{ad}}$ is the chosen ET-AUVs for update, and $\text{ran}(X,n)$ selects $n$ experience from $X$ randomly. The set $R^{*}$ denotes all ET-AUVs excluding $AUV_{\text{be}}$.
The specific selection process is as follows:
  1. 1.
    If the current ET-AUV has the highest reward, then randomly select two from the remaining ET-AUVs to avoid the algorithm getting stuck in local optima.
  2. 2.
    If it doesn’t, then only randomly select one from the other ET-AUVs to increase the proportion of experiences from the best ET-AUV.
After obtaining $AUVad$, exclude $AUVad$ from the ET-AUV set to obtain $AUVord$ (Eq. 17).
$AUV_{\text{ord}}=R_{f}\backslash AUV_{\text{ad}}$ (17)
Then, concatenate the state and action information of ET-AUVs in $AUVad$ to form $m1$ (Eq. 18).
$m_{1}=cat_{k\in AUV_{ad}}(o_{k},a_{k})$ (18)
Apply neural networks and dynamic-switching attention weight processing to ET-AUVs in $AUVord$ to form $m2$ (Eq. 19).
$m_{2}=\sum_{k\in AUV_{ord}}\omega_{k}f(o_{k},a_{k})$ (19)
where the dynamic-switching attention weight $\omega_{k}$, is to evaluate the importance of the information of the other AUVs, and $f(o_{k},a_{k})$ is a function to extract key features from $o_{k}$ and $a_{k}$ of the $k$-th ET-AUV.
$\omega_{k}=\frac{r_{fk}}{\sum_{m\in AUV_{ord}}r_{fm}}$ (20)
where $r_{fk}$ is the reward of $k$-th ET-AUV at $f$-th round.
Finally, input $m1$ and $m2$ into a neural network (MLP) to calculate the Q-value function (Eq. 21).
$Q^{\psi}_{i}(o,a)=f_{i}(m_{1},m_{2})$ (21)
where $f_{i}$ is a three-layer MLP.
Dynamic-switching resampling mechanism: In our study, to enhance the training efficiency by using better samples, we propose an improved sampling strategy based on the experience replay buffer $D$. The buffer $D$ accumulates data from multiple training rounds. This data is stored in the form of a series of tuples $D_{i}$, where each tuple records the states, actions, and rewards of all ET-AUVs in a specific round.
Our approach involves extracting a subset $D_{1}$ comprising $n$ samples from $D$ in each training round, which is then used to update the Critic network. Subsequently, the update process selects the top 50% highest-reward samples from $D_{1}$. These selected samples form a new set $D_{2}$, whose samples are then used for a second update for the Critic network. This mechanism ensures that the focus of network learning is on the most valuable experiences.
Network updates: Further, we employ a progressive strategy to update network parameters using the soft update method (Eq. 22), gradually adjusting the target network parameters to stabilize the training process.
$\hat{\omega}\leftarrow\tau\omega+(1-\tau)\hat{\omega}$ (22)
where $\omega$ represents the parameters of the original network and $\bar{\omega}$ represents the parameters of the target work.
Since all Critic networks share common features, we establish a shared loss function to centrally optimize all the Critics (Eq. 23). This loss function (regarded as a measure of training effectiveness) is based on the sum of differences between expected and actual Q-values.
$\begin{array}[]{c}\mathcal{L}_{Q}(\psi)=\sum_{i=1}^{N}E_{\left(o,a,r,o^{\prime}\right)}\left[\left(Q_{i}^{\psi}(o,a)-y_{i}\right)^{2}\right],\end{array}$ (23)
where
$y_{i}=r_{i}+\gamma Q_{i}^{\bar{\psi}}\left(\mathrm{o}^{\prime},a^{\prime}\right)$ (24)
To slightly adjust the policy of each agent, we utilize gradient ascent to update their policy parameters (Eq. 25).
$\nabla_{\theta_{i}}J\left(\pi_{\theta}\right)=E_{o,a\sim\mathcal{D}}\left[\nabla_{\theta_{i}}\pi_{\theta_{i}}\left(a_{i}\mid o_{i}\right)\nabla_{a_{i}}Q_{i}^{\psi}(o,a)\right]$ (25)
where $a_{i}=\pi_{\theta_{i}}(o_{i})$.
V-C Proposed cooperative tracking policy
In this section, we introduce the proposed reward function. In reinforcement learning, by modeling the problem as MDP, the problem can be transformed into a corresponding maximization of the cumulative expected reward. Therefore, we define the reward function as follows (Eq. 26):
$R_{i}(s,a)=\alpha r_{i}^{\delta}+\beta r_{i}^{\zeta}+\delta r_{i}^{\mu}+\gamma r_{i}^{\sigma}$ (26)
where $r_{i}^{\delta}$ is the tracking reward, $r_{i}^{\zeta}$ is the collision-avoidance reward, $r_{i}^{\mu}$ is the ocean current stability reward, and $r_{i}^{\sigma}$ is the velocity matching reward.
For ET-AUV $i$, the tracking reward $r_{i}^{\delta}$ is designed to control the distance between EC-AUV and the tracking target, thereby maintaining the agent and the target within an best tracking distance range (Eq. 27).
$r_{i}^{\delta}=-\omega_{1}\frac{|d^{\phi}_{(i,j)}-d^{\phi}_{\text{be}}|}{d^{\phi}_{\text{be}}}$ (27)
where $d^{\phi}_{(i,j)}$ is the distance between ET-AUV $i$ and target $j$, and $d^{\phi}_{\text{be}}$ is the best tracking distance.
At the same time, to ensure a safe distance between ET-AUVs during the tracking process, we introduce a collision-avoidance reward $r_{i}^{\zeta}$ (Eq. 28):
$r_{i}^{\zeta}=-\omega_{2}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\frac{|d^{\psi}_{(i,j)}-d^{\psi}_{\text{be}}|}{d^{\psi}_{\text{be}}}$ (28)
where $d^{\psi}_{(i,j)}$ is the distance between ET-AUV $i$ and $j$, and $d^{\psi}_{\text{be}}$ is the best distance between ET-AUVs.
To promote the smooth movement of ET-AUVs under the influence of ocean currents, we introduce an ocean current stability reward $r_{i}^{\mu}$ (Eq. 29). This reward helps mitigate the effects on ET-AUV’s velocity caused by ocean currents.
$r_{i}^{\mu}=-\omega_{3}\frac{\lVert v_{\text{cur}}(i)-v_{\text{pre}}(i)\rVert}{\lVert v_{\text{pre}}(i)\rVert}$ (29)
where $v_{\text{cur}}(i)$ is ET-AUV $i$’s current velocity, $v_{\text{pre}}(i)$ is ET-AUV $i$’s previous velocity.
To avoid excessive energy consumption by ET-AUVs due to frequent acceleration and deceleration, we introduce a velocity matching reward $r_{i}^{\sigma}$ (Eq. 30):
$r_{i}^{\sigma}=-\omega_{4}\frac{\lVert v_{\text{age}}(i)-v_{\text{tar}}(j)\rVert}{\lVert v_{\text{tar}}(j)\rVert}$ (30)
where $v_{\text{age}}(i)$ is the velocity of the agent, $v_{\text{tar}}(j)$ is the velocity of the target.
Based on the AUV swarm network, the cooperative underwater target tracking algorithm can be summarized as shown in Algorithm 3.
The proposed Algorithm 3 combines the proposed DSBM architecture with the proposed tracking policy. This integration enables accurate and cooperative performance in underwater cooperative target tracking, while the ET-AUV’s endurance and resistance to ocean current interference is taken into account.
Algorithm 3 FATHOM-Net-based DSBM tracking algorithm
1:Initialize the Set $N_{E}$, $T$, $N$, respectively.
2:Initialize the weights of Actor and Critic networks corresponding to each ET-AUV.
3:Initialize a random process $\mathcal{N}$ for exploration of action.
4:Initialize replay buffer, environment.
5:for $\text{Episode}=1$ to $N_{E}$ do
6:     Reset environment.
7:     Assign tracking formations according to Algorithm 2.
8:     for $t=1$ to $T$ do
9:         Each ET-AUV takes action $a_{i}=\pi_{\theta_{i}}(o_{i})+\mathcal{N}_{t}$ based on the policy and exploration.
10:         Calculate current force according to Algorithm 1.
11:         Update environment, obtain next state and reward.
12:         Add experience to the replay buffer.
13:         for $i=1$ to $N$ do
14:              Select ET-AUVs based on dynamic-switching attention mechanism.
15:              Calculate $Q_{i}(\psi)$ based on selected ET-AUVs.
16:              Update the actor according to Eq. 25.
17:              Update the Critic according to Eq. 23.
18:     Perform dynamic-switching resampling mechanism from $(x_{t},a_{t},r_{t},x_{t+1})$.
19:     Update the target network of each ET-AUV according to Eq. 22.
VI Evaluations
This section presents the evaluation results, where we conduct an in-depth analysis of the newly proposed tracking algorithm from two perspectives: the effectiveness of the algorithm and the performance of target tracking. We demonstrate our proposed algorithm by comparing it with some of current mainstream MARL algorithms.
VI-A Simulation setup
The test is conducted on a computer equipped with an Intel(R) Core(TM) i9-12900H 2.50GHz processor and 32GB of RAM.
We construct a model using coordinate gridization, treating ET-AUVs as moving particles in a 3D underwater environment. The targets initially locate near the origin and move with a random velocity, while the ET-AUVs are uniformly distributed within a region one kilometer away from the targets.
The relevant evaluation notations is detailed in Table II.
TABLE II: Simulation notation
Notation Description Value
$N_{A}$ Number of ET-AUVs [4,6,12]
$N_{T}$ Number of Targets [2,3,4]
$l_{r}$ Learning rate $33e-$ 3
$N_{E}$ Number of training rounds 5000
$N_{h}$ Number of neurons in the hidden layer 64
$\gamma$ Discount factor 0.95
$\tau$ Network update coefficient $21e-$ 2
$d^{\phi}_{\text{min}}$ Minimum tracking distance 80 m
$d^{\kappa}_{\text{min}}$ Minimum distance among ET-AUVs 80 m
Figure 4: Convergence speed comparison under stable ocean currents (4 ET-AUVs tracking 2 targets)
Figure 5: Convergence speed comparison under ocean current interference (4 ET-AUVs tracking 2 targets)
Figure 6: Convergence speed comparison under stable ocean currents (12 ET-AUVs tracking 4 targets)
Figure 7: Convergence speed comparison under ocean current interference (12 ET-AUVs tracking 4 targets)
VI-B Results
First, to measure the effectiveness of the proposed DSBM, we compare DSBM, with MADDPG [36], MAAC [37], and MADDPG-SAC [38]. Note that MADDPG-SAC is based on MADDPG where the entropy value (that is adaptively adjusted by SAC), is incorporated into the optimization goal to change the update of the Critic network. We compare DSBM with MADDPG-SAC to evaluate the significance of entropy value in the algorithm comparison. We conduct the comparison by respectively utilizing different algorithms to tracking the underwater targets in terms of convergence speed, tracking distance maintenance, difference in velocity from target, and intra-formation strategy consistency.
Convergence speed: To evaluate the stability of algorithm convergence, we repeat multiple rounds of experiments (under different tracking scenarios) and select the samples in a 95% confidence interval. Totally, as shown in Fig. 4 to Fig. 7, our proposed DBSM performs the best than the other three compared objects in convergence speed.
Figure 8: Distribution of Agent-Target distance with stable ocean currents
Figure 9: Distribution of Agent-Target distance with ocean current interference
Figure 10: Agent-Target velocity difference with stable ocean currents
Figure 11: Agent-Target velocity difference with ocean current interference
In details, in Fig. 4, the proposed DSBM showcases the best convergence performance, even if the other algorithms eventually show a trend of convergence. It’s worth noting, MADDPG-SAC converges the fastest in the early stages of training, but the final convergence effect is not satisfactory. This is because the entropy value-based mechanism in MADDPG-SAC improves the exploration scope in the early stage of training, but it also results in that the training to be unstable in the later stages. Different with the case, as shown in Fig. 5, where the ocean current interference is employed, only the DSBM algorithm is able to converge. This demonstrates that our proposed DSBM’s superior adaptability to ocean currents.
Figure 12: Strategic consistency with stable ocean currents
Figure 13: Strategic consistency with ocean current interference
Furthermore, as shown in Fig. 6 and Fig. 7, DSBM also showcases the fastest convergence speed and the best performance with/without ocean current interference.
The results in Fig. 4, Fig. 5, Fig. 6 and Fig. 7 demonstrate that: DSBM pays more attention to the information of better-performing ET-AUVs and conducts more meaningful exploration, thus the time the ineffective exploration can be significantly reduced.
Tracking distance maintenance: As it is aforementioned, maintaining the distance between the agent and the target, i.e., the ”Agent-Target distance” during the tracking process is crucial. Thus, we test as well the distribution of Agent-Target distance with/without ocean current interference.
As the results shown in Fig. 8, with stable ocean current, DSBM can maintain the distance between ET-AUV and its targets within a small (pre-defined) distance change range. In contrast, the other algorithms all display significant deviations. Further, in Fig. 9, the similar results are also achieved in the scenario with ocean current interference.
Velocity difference maintenance: In AUV swarm network-based underwater cooperative tracking, frequent acceleration and deceleration for ET-AUV can rapidly consume the energy, making against the long-duration tracking tasks especially in underwater. Therefore, the velocity difference between the ET-AUV and target, i.e., the ”Agent-Target velocity difference” can be used to measure the endurance capacity and tracking efficiency of the tracking policy.
As shown in Fig. 10, with stable ocean current conditions, the Agent-Target velocity differences of DSBM are within a very small range. In contrast, the other three algorithms showcase a wider range of Agent-Target velocity differences, indicating that the ET-AUVs have to frequently accelerate and decelerate throughout the entire tracking process. This will significantly reduce the endurance time under the same state conditions.
As shown in Fig. 11, with ocean current interference, the velocity of ET-AUVs guided by DSBM almost maintains consistency with the velocity of the tracking target. The results in both Fig. 10 and Fig. 11 demonstrate that the proposed DSBM can perform stable and high-efficient underwater cooperative tracking with complex underwater environment (with/without ocean current interference).
(a) Scenario 1_Phase 1
(b) Scenario 1_Phase 2
(c) Scenario 1_Phase 3
(d) Scenario 2_Phase 1
(e) Scenario 2_Phase 2
(f) Scenario 2_Phase 3
(g) Scenario 3_Phase 1
(h) Scenario 3_Phase 2
(i) Scenario 3_Phase 3
Figure 14: Test for the availability of the proposed approaches in 3D environment.
Intra-formation strategy consistency: During the tracking process, a consistent strategy for the AUV formation helps improve efficiency in target identification and tracking, especially in complex underwater environments. This is crucial to prevent tracking errors or relief target loss due to environmental factors. Hence, we take the experimental scenario where twelve ET-AUVs tracking four targets as an example. And the ET-AUVs in each formation each choose actions from seven possible actions, to evaluate the consistency of these actions.
As the results shown in both Fig. 12 and 13, we showcase the probability distribution of the number of the consistent ET-AUVs in the formation. As shown in Fig. 12, before the ocean current interference is taken into account, the actions of all the ET-AUVs under our proposed DSBM are more likely consistent. Similarly, the proposed DSBM obviously performs the best than the other compared objects, while the ocean current interference is considered. This benefits from the proposed DSBM is based on our proposed FATHOM-Net framework, by which the AUV formation within a region undergoes local training, focusing more on the tracking targets within the region rather than the entire tracking task.
Finally, to make the entire tracking process directed by our proposed DSBM more clear, we use Unity to simulate the underwater environment for visualizing the tracking process. As shown in Fig. 14, the cubes represent the targets, and the spheres represent ET-AUVs. The red solid line represents the movement trajectory of the target, while other colored solid lines represent the tracking trajectories of the ET-AUVs. Specifically, we simulate three scenarios of multi-target tracking. Fig. 14(a)-Fig. 14(c) represent four ET-AUVs tracking two targets, Fig. 14(d)-Fig. 14(f) represent six ET-AUVs tracking three targets, and Fig. 14(g)-Fig. 14(i) represent twelve ET-AUVs tracking four targets. The visualization in Fig. 14 also proves the effectiveness of the proposed DSBM in tracking underwater targets in a 3D environment.
VII Conclusion
In this study, we propose a dynamic-switching-enabled multi-agent reinforcement learning framework for performing AUV swarm-network-based cooperative multi-target tracking in underwater environment. We first give modelling approach for the underwater sonar-based detection and the ocean current interference. Then, we propose FATHOM-Net, a novel MARL architecture based on SDN, which enhances the flexibility and robustness of the AUV swarm. The proposed DSBM includes the ”dynamic-switching attention” and ”dynamic-switching resampling” mechanisms, both based on an experience replay buffer. It enhances AUV swarm network’s self-learning efficiency and accuracy. Finally, to achieve efficient formation classification in the interference of ocean currents, we propose an algorithm named ASMA. It efficiently and accurately assigns formations for ET-AUV through fuzzy logic and rule-based expert systems. Evaluation results demonstrate that our proposed tracking algorithm achieves the fastest convergence speed and tracking performance compared to various popular research products. Future research directions derived from this work can be summarized as the following:
  1. 1.
    Optimizing underwater obstacle avoidance mechanisms for ET-AUVs to mitigate potential damage;
  2. 2.
    Balancing energy consumption among the AUVs to enhance the endurance of the AUV swarm network-based cooperative tracking systems;
  3. 3.
    Designing AUV swarm robustness control framework when the unstable underwater communication (e.g., the underwater acoustic-based communication) has to been taken into account.